The equation for the reaction of nitrogen and oxygen to form nitrogen oxide is written as
N2(g)+O2(g)→2NO(g)ΔH=+90.2kJ

a)How many kilojoules are required to form 3.40 g of NO

b)What is the equation for the decomposition of NO?

c)How many kilojoules are released when 4.60 g of NO decomposes to N2 and O2?
Answer #1

a)

Molar mass of NO,

MM = 1*MM(N) + 1*MM(O)

= 1*14.01 + 1*16.0

= 30.01 g/mol

mass(NO)= 3.40 g

use:

number of mol of NO,

n = mass of NO/molar mass of NO

=(3.4 g)/(30.01 g/mol)

= 0.1133 mol

Since Δ H is positive, heat is absorbed

when 2 mol of NO reacts, heat absorbed = 90.2 KJ

So,

for 0.1133 mol of NO, heat absorbed = 0.1133*90.2/2 KJ

= 5.11 KJ

Answer: 5.11 KJ

b)

2NO(g) —> N2(g)+O2(g) ΔH=-90.2kJ

c)

Molar mass of NO,

MM = 1*MM(N) + 1*MM(O)

= 1*14.01 + 1*16.0

= 30.01 g/mol

mass(NO)= 4.60 g

use:

number of mol of NO,

n = mass of NO/molar mass of NO

=(4.6 g)/(30.01 g/mol)

= 0.1533 mol

Since Δ H is negative, heat is released

when 2 mol of NO reacts, heat released = 90.2 KJ

So,

for 0.1533 mol of NO, heat released = 0.1533*90.2/2 KJ

= 6.913 KJ

Answer: 6.91 KJ